# Combining Teacher-Student with Representation Learning: A Concurrent Teacher-Student Reinforcement Learning Paradigm for Legged Locomotion

### v1

@@ Abstract
Thanks to the explosive developments of data-driven learning methodologies recently, reinforcement learning (RL) emerges as a promising solution to address the legged locomotion problem in robotics. In this manuscript, we propose a novel concurrent teacher-student reinforcement learning architecture for legged locomotion over challenging terrains, based only on proprioceptive measurements in real-world deployment. Different from convectional teacher-student architecture that trains the teacher policy via RL and transfers the knowledge to the student policy through supervised learning, our proposed architecture trains teacher and student policy networks concurrently under the reinforcement learning paradigm. To achieve this, we develop a new training scheme based on conventional proximal policy gradient (PPO) method to accommodate the interaction between teacher policy network and student policy network. The effectiveness of the proposed architecture as well as the new training scheme is demonstrated through extensive indoor and outdoor experiments on quadrupedal robots and point-foot bipedal robot, showcasing robust locomotion over challenging terrains and improved performance compared to two-stage training methods.
@@ I Introduction
Locomotion is one of the most important skills for legged robots, which enables them to traverse complicated terrains to accomplish various tasks [1, 2, 3, 4, 5, 6, 7]. Due to the complicated contact interaction with the uneven terrain and the inherent nonlinear and hybrid dynamics for legged locomotion, the locomotion controller synthesis problem is widely acknowledged to be challenging [8]. Recently, reinforcement learning-based methods have been shown to be a promising solution to legged locomotion and have achieved remarkable results [9, 10, 11, 12, 13].
Nowadays, the so-called teacher-student paradigm is one of the most widely adopted and studied methodology for achieving legged locomotion with reinforcement learning [14, 10, 15]. In such a paradigm, a teacher policy having full access to all information (privileged) related to locomotion (e.g., terrain details, contact information, accurate inertial parameters, etc.) is trained by reinforcement learning. Then, a student policy that operates purely with proprioceptive feedback is trained by supervised learning to imitate the behavior of trained teacher policy. This approach allows for efficient learning and sim-to-real transfer of the policy to real-world where the robot operates with only proprioception. Such a methodology has been widely extended and been shown effective in enabling various legged robots to achieve locomotion in complex terrains [10, 15, 16].  [15] integrates teacher-student learning with Adversarial Motion Priors (AMP) to enable quadruped to learn natural and robust locomotion.  [16] leverages the advances in rapid motor adaptation for quadruped locomotion proposed originally in [10], and expands them to humanoid robots. Similar training paradigm has also been shown capable of addressing other challenging locomotion tasks with richer inputs including exteroceptive measurements [17, 18, 19, 20, 21].
@@
Figure 1: Our training method enables legged robots of various sizes and configurations to achieve robust and agile locomotion across challenging real-world terrains, while also possessing exceptional capabilities to withstand strong external disturbances.
Different from the teacher-student paradigm, representation learning integrates dynamics learning to enhance legged locomotion [22, 23]. DreamWaQ [22] leverages learned representations through variational autoencoders (VAE) to improve legged locomotion performance. Hybrid Internal Model [23] treats external states like terrain conditions as disturbances. It learns representations through contrastive learning, which pulls close the representations generated from historical sequence of observations and future observations.A common characteristic of these methodologies is the use of a regression objective for state representation, which requires the neural network to match the targets accurately.
Most of the aforementioned RL-based strategies are validated with quadruped platforms over challenging terrains. Extensions of these methods to bipedal robots have not been fully investigated.  [24] achieved the sim-to-real transfer of a RNN policy for bipedal walking on flat ground.  [25] employs a terrain randomization approach to implement proprioceptive controllers for bipedal robots involved in climbing and descending stairs. However, these methods employ vanilla reinforcement learning to train policy networks while overlooking the full utilization of privileged information, suggesting that there is considerable room to enhance the blind locomotion capabilities of bipedal robots.
In this work, we explore the mechanism for systematically integrating the information from the teacher strategy with that from the student strategy. To this end, we propose a concurrent teacher-student learning architecture for legged locomotion, which philosophically combines the teacher-student paradigm with the core of representation learning and embodies the training of teacher and student policies concurrently. Based on teacher-student paradigm, we develop a variant of the PPO training algorithm to incorporate the iterative training of the teacher and student policy to ensure convergence of the overall training process. We validate our approach with multiple hardware platforms including quadrupeds of different sizes and with a more challenging, underactuated bipedal robot with point feet. Furthermore, robots have demonstrated the ability to navigate through challenging terrains and against external disturbances.
@@
Figure 2: Overview of the learning framework. The teacher and student policies are trained concurrently using PPO within an asymmetric actor-critic framework. Agents in both groups share the same policy and critic networks, with actions determined by observations and latent representations from either privileged or proprioceptive encoder. The privileged encoder is trained via policy gradient, while the proprioceptive encoder undergoes supervised learning to minimize reconstruction loss.
@@
The contributions of this work are summarized as follows. First, we develop a novel concurrent teacher-student reinforcement learning architecture that integrates standard teacher-student paradigm and representation learning to exploit the interplay between teacher and student networks to enhance the reinforcement learning performance. Second, we show that the proposed architecture exhibits significant generalizability and adaptability across different locomotion problems with different hardware platforms with minimal modifications. Third, extensive hardware experimental results with various hardware platforms, including quadrupeds and point-foot biped, are demonstrated, achieving robust locomotion over challenging terrains in both indoor and outdoor environments.
@@ II-A Legged Locomotion Problem and Reinforcement Learning
In essence, legged locomotion focuses on finding the appropriate joint torque commands for all actuated joints of the robots given the sensory measurements. Under the assumption of only having accessibility to priopricetive measurements from IMU and joint encoders, the legged locomotion dynamics can be formulated as the following infinite-horizon partially observable Markov decision process (POMDP), defined by the tuple $\mathcal{S},\mathcal{A},\mathcal{O},T,\Omega,R\right>$, where $\mathcal{S}\subset\mathbb{R}^{n}$ is the set of full state including all dynamic information of legged robot and environment around, $\mathcal{A}\subset\mathbb{R}^{m}$ is the set of action, $\mathcal{O}\subset\mathbb{R}^{o}$ is the set of observation, $T(\boldsymbol{s}^{\prime},\boldsymbol{s},\boldsymbol{a})=p(\boldsymbol{s}^{\prime}|\boldsymbol{s},\boldsymbol{a})$ is the state transition function, $\Omega(\boldsymbol{o},\boldsymbol{s},\boldsymbol{a})=p(\boldsymbol{o}|\boldsymbol{s},\boldsymbol{a})$ is the observation function and $R(\boldsymbol{s},\boldsymbol{a},\boldsymbol{s}^{\prime})$ is the reward function. Our goal is to find the optimal policy $\pi^{*}$ to maximize the expected discounted return over the trajectory:
@@
State Space: We denote by $\boldsymbol{o}_{t}$ the observation and $\boldsymbol{s}_{t}$ the state at time $t$. The proprioceptive observation $\boldsymbol{o}_{t}\in\mathbb{R}$ consists of angular velocity, gravity vector in the base frame of the robot, joint positions and velocities, command, and previous actions. The command consists of the desired velocity $v^{\text{cmd}}_{x},v^{\text{cmd}}_{y}$ and the angular velocity $\omega^{\text{cmd}}_{z}$ in the base frame. The full state $\boldsymbol{s}_{t}$ consists of proprioceptive observation $\boldsymbol{o}_{t}$, base linear velocity $\boldsymbol{v}_{t}\in\mathbb{R}^{3}$, terrain height samples $i_{t}\in\mathbb{R}$ and other privileged information including feet contact forces, joint torques and acceleration of joints.
Action Space: For each actuated joint, the action $\boldsymbol{a}_{t}\in\mathbb{R}$ represents the angular deviation of the robot’s joint relative to its nominal position. Hence, the robot’s joint PD controller reference is
@@
II-B Concurrent Teacher-Student Network Architectures
@@
We train the teacher and student policy concurrently by dividing parallel agents into two groups named teacher group and student group, then employing the asymmetric actor-critic framework shown in Fig. 2. Agents in two groups are both trained using proximal policy optimization (PPO) while they share the same policy network $\pi_{\theta}$ and critic network $V_{\phi}$ designed as the Multi-Layer Perceptron (MLP) with Exponential Linear Unit (ELU) activation. The policy network outputs action $\boldsymbol{a}_{t}$ given proprioceptive observation $\boldsymbol{o}_{t}$ and latent representation $\boldsymbol{z}_{t}\in\mathbb{R}^{32}$, which were generated by the precedent encoder designed as MLP with ELU activation. The latent representation $\boldsymbol{z}_{t}$ is mapped onto a unit hypersphere with a $\mathcal{L}_{2}$-normalization. The critic network estimates the state value given state $\boldsymbol{s}_{t}$ and latent representation $\boldsymbol{z}_{t}$.
Throughout this manuscript, we use superscripts $(\cdot)^{\text{t}}$ and $(\cdot)^{\text{s}}$ to indicate teacher group and student group, respectively. Agents in teacher group have access to full state $\boldsymbol{s}_{t}$ while they utilize privileged encoder $E_{\theta^{\text{t}}}$ to encode state $\boldsymbol{s}^{\text{t}}_{t}$ into latent representation $\boldsymbol{z}^{\text{t}}_{t}$. Agents in the student group only have access to proprioceptive observation, allowing policies to be deployed in real environments. Proprioceptive encoder $E_{\theta^{\text{s}}}$ is leveraged to encode observation sequence $\boldsymbol{o}^{\text{s}}_{t-H:t}=[\boldsymbol{o}^{\text{s}}_{t},\cdots,\boldsymbol{o}^{\text{s}}_{t-H}]^{T}$ into latent representation $\boldsymbol{z}^{\text{s}}_{t}$ similar to privileged encoder.
The latent representation $\boldsymbol{z}_{t}$ is generated from the privileged encoder or the proprioceptive generator, which guides the policy network to produce specific actions for various terrains and scenarios. The privileged encoder and policy network is trained through policy gradient to maximize the expected discounted return, while the proprioceptive encoder is trained through supervised learning to minimize the reconstruction loss between outputs of proprioceptive encoder and privileged encoder. More details on each network are shown in Table I.
@@
Module Notation Inputs Outputs
Privileged Encoder $E_{\theta^{\text{t}}}$ $\boldsymbol{s}^{\text{t}}_{t}$ $\boldsymbol{z}^{\text{t}}_{t}$
Proprioceptive Encoder $E_{\theta^{\text{s}}}$ $\boldsymbol{o}^{\text{s}}_{t-H:t}$ $\boldsymbol{z}^{\text{s}}_{t}$
Policy Network $\pi_{\theta}$ $\boldsymbol{o}_{t},\boldsymbol{z}_{t}$ $\boldsymbol{a}_{t}$
Critic $V_{\phi}$ $\boldsymbol{s}_{t},\boldsymbol{z}_{t}$ $\hat{V}_{t}$
@@ II-C Training Pipeline
where $\mathcal{D}^{\text{t}}$ and $\mathcal{D}^{\text{s}}$ are sets of teacher group and student group trajectories by interacting with environment using $E_{\theta^{\text{t}}_{k}},\pi_{\theta_{k}}$ and $E_{\theta^{\text{s}}_{k}},\pi_{\theta_{k}}$, respectively. $T$ is the length of corresponding trajectory. $r^{\text{t}}_{t},r^{\text{s}}_{t}$ are ratio functions of two groups:
$r^{\text{t}}_{t}(\theta,\theta^{\text{t}})=\frac{\pi_{\theta}\left(\boldsymbol{a}^{\text{t}}_{t}|\boldsymbol{o}^{\text{t}}_{t},E_{\theta^{\text{t}}}(\boldsymbol{s}^{\text{t}}_{t})\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{a}^{\text{t}}_{t}|\boldsymbol{o}^{\text{t}}_{t},E_{\theta_{\text{old}}^{\text{t}}}(\boldsymbol{s}^{\text{t}}_{t})\right)}$ (5)
$r^{\text{s}}_{t}(\theta)=\frac{\pi_{\theta}\left(\boldsymbol{a}^{\text{s}}_{t}|\boldsymbol{o}^{\text{s}}_{t},E_{\theta^{\text{s}}}(\boldsymbol{o}_{t-H:t}^{\text{s}})\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{a}^{\text{s}}_{t}|\boldsymbol{o}^{\text{s}}_{t},E_{\theta^{\text{s}}}(\boldsymbol{o}_{t-H:t}^{\text{s}})\right)}$ (6)
Value function $V_{\phi}$ is trained by regression on mean-square error between $V_{\phi}(\boldsymbol{s}_{t},\boldsymbol{c}_{t},\boldsymbol{z}_{t})$ and $\hat{R}_{t}$ estimated by Generalized Advantage Estimation (GAE) using trajectories from both groups. The Monte-Carlo approximation of value loss is defined as:
@@
$L^{\text{rec}}(\theta^{\text{s}})=\frac{1}{|\mathcal{D}^{\text{s}}|T}\sum_{\tau\in\mathcal{D}^{\text{s}}}\sum_{t=0}^{T}\left\|E_{\theta^{\text{s}}}(\boldsymbol{o}_{t-H:t}^{\text{s}})-E_{\theta^{\text{t}}}(\boldsymbol{s}^{\text{s}}_{t})\right\|_{2}^{2}$ (8)
@@
3:     Collect sets of trajectories $\mathcal{D}_{k}$ with latest policy

### v2

# CTS: Concurrent Teacher-Student Reinforcement Learning for Legged Locomotion

@@ Abstract
Thanks to recent explosive developments of data-driven learning methodologies, reinforcement learning (RL) emerges as a promising solution to address the legged locomotion problem in robotics. In this paper, we propose CTS, a novel Concurrent Teacher-Student reinforcement learning architecture for legged locomotion over uneven terrains. Different from conventional teacher-student architecture that trains the teacher policy via RL first and then transfers the knowledge to the student policy through supervised learning, our proposed architecture trains teacher and student policy networks concurrently under the reinforcement learning paradigm. To this end, we develop a new training scheme based on a modified proximal policy gradient (PPO) method that exploits data samples collected from the interactions between both the teacher and the student policies with the environment. The effectiveness of the proposed architecture and the new training scheme is demonstrated through substantial quantitative simulation comparisons with the state-of-the-art approaches and extensive indoor and outdoor experiments with quadrupedal and point-foot bipedal robot platforms, showcasing robust and agile locomotion capability. Quantitative simulation comparisons show that our approach reduces the average velocity tracking error by up to 20% compared to the two-stage teacher-student, demonstrating significant superiority in addressing blind locomotion tasks. Videos are available at https://clearlab-sustech.github.io/concurrentTS/.
Index Terms:
Legged Robots, Reinforcement Learning, Machine Learning for Robot Control
@@ I Introduction
Lcomotion is one of the most important skills for legged robots, which enables them to traverse complicated terrains to accomplish various tasks [1, 2, 3, 4, 5, 6, 7]. Due to the complicated contact interaction with the uneven terrain and the inherent nonlinear and hybrid dynamics, the locomotion controller synthesis problem is widely acknowledged to be challenging [8]. Recently, reinforcement learning-based methods have been shown to be a promising solution to legged locomotion and have achieved remarkable results  [9, 10, 11, 12, 13, 14, 15].
@@
Figure 1: CTS enables legged robots of various sizes and configurations to achieve robust and agile locomotion across challenging real-world terrains, while also possessing exceptional capabilities to withstand strong external disturbances.
@@
Figure 2: Overview of the learning framework. The teacher and student policies are trained concurrently using PPO within an asymmetric actor-critic framework. Agents in both groups share the same critic and policy network, with actions determined by observations and latent representations from either privileged or proprioceptive encoder. The privileged encoder is trained via policy gradient, while the proprioceptive encoder undergoes supervised learning to minimize reconstruction loss.
Nowadays, the teacher-student paradigm is one of the most widely adopted and studied learning-based methodologies for achieving legged locomotion [16, 12, 17]. In such a paradigm, a teacher policy having full access to all locomotion-related (privileged) information (e.g., terrain details, contact information, accurate inertial parameters) is first trained by reinforcement learning. Then, a student policy that operates purely with proprioceptive feedback is trained by supervised learning to reconstruct the latent representation from the encoder of trained teacher policy and/or imitate the actor output of teacher policy. This approach allows for efficient learning and sim-to-real transfer where the robot operates with only proprioception, enabling various legged robots to locomote in complex terrains [12, 17, 18]. [17] integrates teacher-student learning with Adversarial Motion Priors (AMP) to enable quadruped robots to learn natural and robust locomotion.  [18] leverages the advances in rapid motor adaptation for quadruped locomotion originally proposed in [12], and extends the application to humanoid robots. Similar training paradigm has also been shown capable of addressing other challenging locomotion tasks with richer inputs including exteroceptive measurements [19, 20, 21, 22, 23]. In addition to the above two-stage training process, the Regularized Online Adaptation (ROA) method, as proposed in [24] integrates the training of encoders of the teacher and student policies into a single stage, achieving mobile-manipulation tasks. During the iterations in ROA where the policy uses the output of the proprioceptive encoder as input, the policy network itself is not updated. In these iterations, only the proprioceptive encoder undergoes supervised learning. This means the policy network does not undergo reinforcement learning training conditioned on the proprioceptive encoder’s input.
Different from the teacher-student paradigm, representation learning leverages the idea of dynamics learning to enhance legged locomotion [25, 26]. DreamWaQ [25] exploits learned representations through variational autoencoders (VAE) to improve legged locomotion performance. Hybrid Internal Model [26] (HIM) treats external states such as terrain conditions as disturbances and learns representations through contrastive learning. In this way, the HIM method closes the gap between the representations generated from historical sequence of observations and future observations, aiding in obtaining latent representations that include system dynamics to assist in reinforcement learning training. A common characteristic of these methodologies is the use of a regression objective for state representation, which requires the neural network to match the targets accurately.
In this work, we explore the mechanism for systematically integrating the information from the teacher strategy with that from the student strategy. To this end, we propose a concurrent teacher-student learning architecture for legged locomotion. Such an architecture philosophically combines the teacher-student paradigm with the core of representation learning and embodies the training of teacher and student policies concurrently, making it more streamlined than two-stage training. More importantly, we guide the student’s training using reinforcement learning objectives instead of merely imitating the teacher. Since the differences in observations between the student and the teacher make perfect imitation difficult, the goal of purely imitating the teacher may not be optimal. Incorporating reinforcement learning objectives during the student’s training helps in deriving a better policy. We validate our approach on multiple hardware platforms including quadrupeds of different sizes and a more challenging, underactuated bipedal robot with point feet. Furthermore, robots have demonstrated the ability to navigate through challenging terrains and against external disturbances.
@@
The contributions of this work are summarized as follows. First, the proposed CTS reinforcement learning architecture effectively exploits the interplay between teacher and student networks to enhance the overall performance of the resultant policy. Simulation experiments quantitatively show that the proposed CTS architecture achieves an improvement of up to 20% in velocity tracking error on uneven terrains compared to state-of-the-art teacher-student and ROA methods. Second, we conducted extensive real-world experiments with various hardware platforms to demonstrate the small sim-to-real gap of the proposed architecture. The extensive hardware experimental results with various hardware platforms, including quadrupeds and point-foot bipeds, illustrate the capability of our proposed method to enable robust locomotion over challenging terrains in both indoor and outdoor environments.
@@ II-A Legged Locomotion Problem and Reinforcement Learning
In essence, legged locomotion focuses on finding the appropriate joint torque commands for all actuated joints of the robots given the sensory measurements. Under the assumption of only having accessibility to proprioceptive measurements from IMU and joint encoders, the legged locomotion dynamics can be formulated as an following infinite-horizon partially observable Markov decision process (POMDP), defined by the tuple $\mathcal{S},\mathcal{A},\mathcal{O},T,\Omega,R\right>$, where $\mathcal{S}\subset\mathbb{R}^{n}$ is the set of full state including all dynamic information of legged robot and environment around, $\mathcal{A}\subset\mathbb{R}^{m}$ is the set of action, $\mathcal{O}\subset\mathbb{R}^{o}$ is the set of observation, $T(\boldsymbol{s}^{\prime},\boldsymbol{s},\boldsymbol{a})=p(\boldsymbol{s}^{\prime}|\boldsymbol{s},\boldsymbol{a})$ is the state transition function, $\Omega(\boldsymbol{o},\boldsymbol{s},\boldsymbol{a})=p(\boldsymbol{o}|\boldsymbol{s},\boldsymbol{a})$ is the observation function and $R(\boldsymbol{s},\boldsymbol{a},\boldsymbol{s}^{\prime})$ is the reward function. Our goal is to find the optimal policy $\pi^{*}$ to maximize the expected discounted return over the trajectory:
@@
State Space: We denote by $\boldsymbol{o}_{t}$ the observation and $\boldsymbol{s}_{t}$ the state at time $t$. The proprioceptive observation $\boldsymbol{o}_{t}\in\mathbb{R}^{n}$ consists of angular velocity, gravity vector in the base frame of the robot, joint positions and velocities, command, and previous actions. Here, $n$ denotes the dimensionality of the proprioceptive observation vector, encompassing all aforementioned components. The command consists of the desired velocity $v^{\text{cmd}}_{x},v^{\text{cmd}}_{y}$ and the angular velocity $\omega^{\text{cmd}}_{z}$ in the base frame. The full state $\boldsymbol{s}_{t}$ consists of proprioceptive observation $\boldsymbol{o}_{t}$, base linear velocity $\boldsymbol{v}_{t}\in\mathbb{R}^{3}$, terrain height samples $i_{t}\in\mathbb{R}^{m}$ and other privileged information including feet contact forces, joint torques and acceleration of joints. In this context, $m$ represents the number of terrain height samples collected, forming a vector that describes the terrain profile.
Action Space: For each actuated joint, the action $\boldsymbol{a}_{t}\in\mathbb{R}^{k}$ represents the angular deviation of the robot’s joint relative to its nominal position , where $k$ denotes the number of actuated joints. Hence, the robot’s joint PD controller reference is
@@
II-B Concurrent Teacher-Student Architectures
@@
We train the teacher and student policies concurrently by dividing parallel agents into two groups named teacher group and student group, then employing the asymmetric actor-critic framework shown in Fig. 2, where the teacher policy includes privileged encoder and policy network, and the student policy includes proprioceptive encoder and policy network. Agents in two groups are both trained using proximal policy optimization (PPO) while they share the same policy network $\pi_{\theta}$ and critic network $V_{\phi}$. The policy network outputs action $\boldsymbol{a}_{t}$ given proprioceptive observation $\boldsymbol{o}_{t}$ and latent representation $\boldsymbol{z}_{t}\in\mathbb{R}^{32}$, which were generated by the precedent encoder. The latent representation $\boldsymbol{z}_{t}$ is mapped onto a unit hypersphere with a $\mathcal{L}_{2}$-normalization. The critic network estimates the state value for some given state $\boldsymbol{s}_{t}$ and latent representation $\boldsymbol{z}_{t}$.
Throughout this manuscript, we use superscripts $(\cdot)^{\text{t}}$ and $(\cdot)^{\text{s}}$ to indicate teacher group and student group, respectively. Agents in teacher group have access to full state $\boldsymbol{s}_{t}$ while they utilize privileged encoder $E^{\text{t}}_{\theta}$ to encode state $\boldsymbol{s}^{\text{t}}_{t}$ into latent representation $\boldsymbol{z}^{\text{t}}_{t}$. Agents in the student group only have access to proprioceptive observation, allowing policies to be deployed in real environments. Proprioceptive encoder $E^{\text{s}}_{\theta}$ is leveraged to encode observation sequence $\boldsymbol{o}^{\text{s}}_{t-H:t}=[\boldsymbol{o}^{\text{s}}_{t},\cdots,\boldsymbol{o}^{\text{s}}_{t-H}]^{T}$ into latent representation $\boldsymbol{z}^{\text{s}}_{t}$ similar to privileged encoder.
The latent representation $\boldsymbol{z}_{t}$ is generated from the privileged encoder or the proprioceptive encoder, which guides the policy network to produce specific actions for various terrains and scenarios. The privileged encoder and policy network is trained through policy gradient to maximize the expected discounted return, while the proprioceptive encoder is trained through supervised learning to minimize the reconstruction loss between outputs of proprioceptive encoder and privileged encoder. All modules are designed as the Multi-Layer Perceptron (MLP) with Exponential Linear Unit (ELU) activation. More details on each network are shown in Table I.
@@
Module Notation Inputs Hidden Layer Outputs
Privileged Enc. $E^{\text{t}}_{\theta}$ $\boldsymbol{s}^{\text{t}}_{t}$ [512, 256] $\boldsymbol{z}^{\text{t}}_{t}$
Proprioceptive Enc. $E^{\text{s}}_{\theta}$ $\boldsymbol{o}^{\text{s}}_{t-H:t}$ [512, 256] $\boldsymbol{z}^{\text{s}}_{t}$
Policy Network $\pi_{\theta}$ $\boldsymbol{o}_{t},\boldsymbol{z}_{t}$ [512, 256, 128] $\boldsymbol{a}_{t}$
Critic $V_{\phi}$ $\boldsymbol{s}_{t},\boldsymbol{z}_{t}$ [512, 256, 128] $\hat{V}_{t}$
@@ II-C Training Pipeline
where $\mathcal{D}^{\text{t}}$ and $\mathcal{D}^{\text{s}}$ are sets of teacher group and student group trajectories by interacting with environment using $E^{\text{t}}_{\theta_{k}},\pi_{\theta_{k}}$ and $E^{\text{s}}_{\theta_{k}},\pi_{\theta_{k}}$, respectively. $T$ is the length of corresponding trajectory. $r^{\text{t}}_{t},r^{\text{s}}_{t}$ are ratio functions of two groups:
$r^{\text{t}}_{t}(\theta,\theta^{\text{t}})=\frac{\pi_{\theta}\left(\boldsymbol{a}^{\text{t}}_{t}|\boldsymbol{o}^{\text{t}}_{t},E^{\text{t}}_{\theta}(\boldsymbol{s}^{\text{t}}_{t})\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{a}^{\text{t}}_{t}|\boldsymbol{o}^{\text{t}}_{t},E^{\text{t}}_{\theta_{\text{old}}}(\boldsymbol{s}^{\text{t}}_{t})\right)}$ (5)
$r^{\text{s}}_{t}(\theta)=\frac{\pi_{\theta}\left(\boldsymbol{a}^{\text{s}}_{t}|\boldsymbol{o}^{\text{s}}_{t},E^{\text{s}}_{\theta}(\boldsymbol{o}_{t-H:t}^{\text{s}})\right)}{\pi_{\theta_{\text{old}}}\left(\boldsymbol{a}^{\text{s}}_{t}|\boldsymbol{o}^{\text{s}}_{t},E^{\text{s}}_{\theta}(\boldsymbol{o}_{t-H:t}^{\text{s}})\right)}$ (6)
Value function $V_{\phi}$ is trained by regression on mean-square error between $V_{\phi}(\boldsymbol{s}_{t},\boldsymbol{z}_{t})$ and $\hat{R}_{t}$ estimated by Generalized Advantage Estimation (GAE) using trajectories from both groups. The Monte-Carlo approximation of value loss is defined as:
@@
$L^{\text{rec}}(\theta^{\text{s}})=\frac{1}{|\mathcal{D}^{\text{s}}|T}\sum_{\tau\in\mathcal{D}^{\text{s}}}\sum_{t=0}^{T}\left\|E^{\text{s}}_{\theta}(\boldsymbol{o}_{t-H:t}^{\text{s}})-E^{\text{t}}_{\theta}(\boldsymbol{s}^{\text{t}}_{t})\right\|_{2}^{2}$ (8)
Remark 1.
Different from two-stage teacher-student paradigm in which the student policy aims to imitate a well-trained teacher policy, the proposed CTS extensively exploit the interplay between the teacher and the student policies. Such an online iterative training scheme may seem to cause the student to learn a moving target, but this change is gradual, ultimately leading to better student policy performance as shown later in the results section.
@@
3:     Collect sets of trajectories $\mathcal{D}^{\text{t}}$ and $\mathcal{D}^{\text{s}}$ with latest policy